# Towards Consistent and Controllable Image Synthesis for Face Editing

### v1

@@ IV-C Ablation Study
Disentangled and pixel-aligned image condition. To demonstrate the effectiveness of using disentangled conditions and of using image conditions for pose and lighting conditioning, we explore alternative designs, including 1) We concatenate the expression, pose, and lighting coefficients together into a vector, embed this vector through a single linear layer and then add the output into the time embedding of the denoising UNet (No-disent.) 2) We separate the pose, expression and lighting coefficients and deal with them in different linear layers respectively. After that, we concatenate the ouputs from those linear layers into a vector and add it to the time embedding (Coef.-sep.). As shown in the third and fourth rows of Fig. 7, visualizations illustrate that decoupled conditions outperform entangled conditions. Soleing relying on 3DMM coefficients as conditions does not work on pose and lighting editing, which can also cause pose variation when it is not intended to edit. Quantitative results are also presented in Tab. IV, demonstrating the superiority of our design.
Figure 6: Ablation study for different colors used to mask for the background image. Black color is easy to clash with the color of the original images.
@@
Background Color. We find that the color of the masked region has a significant impact on the experimental performance. As demonstrated in Fig. 6, a black background can easily blend with a person’s hair, clothing, or even the original image background, leading to errors in the model’s background identification. This not only affects the preservation of the source image background but also impacts the final editing of the person. To solve this problem, we use gray to mask the human region, as it is a color that rarely clashes with the image background or the person.

### v2

@@ IV-C Ablation Study
Disentangled and pixel-aligned image condition. To demonstrate the effectiveness of using disentangled conditions and of using image conditions for pose and lighting conditioning, we explore alternative designs, including 1) We don’t use image condition. Instead, we concatenate the expression, pose, and lighting coefficients together into a vector to make the conditions entangled. Then we embed this vector through a single linear layer and then add the output into the time embedding of the denoising UNet (No-disent.) 2) We still use the coefficients instead of rendered images for pose and lighting condition, and we separate these coefficients to deal with them in different linear layers respectively. After that, we concatenate the ouputs from those linear layers into a vector and add it to the time embedding (Coef.-sep.). As shown in the third and fourth rows of Fig. 7, visualizations illustrate that decoupled conditions outperform entangled conditions. Soleing relying on 3DMM coefficients as conditions does not work on pose and lighting editing, which can also cause pose variation when it is not intended to edit. Quantitative results are also presented in Tab. IV, demonstrating the superiority of our design.
@@
Foreground Color. We find that the color of the masked region has a significant impact on the experimental performance. As demonstrated in Fig. 6, a black foreground can easily blend with a person’s hair, clothing, or even the original image background, leading to errors in the model’s background identification. This not only affects the preservation of the source image background but also impacts the final editing of the person. To solve this problem, we use gray to mask the human region, as it is a color that rarely clashes with the image background or the person.